### v1

# On the Black-box Explainability of Object Detection Models for Safe and Trustworthy Industrial Applications

Abstract
In the realm of human-machine interaction, artificial intelligence has become a powerful tool for accelerating data modeling tasks. Object detection methods have achieved outstanding results and are widely used in critical domains like autonomous driving and video surveillance. However, their adoption in high-risk applications, where errors may cause severe consequences, remains limited. Explainable Artificial Intelligence (XAI) methods aim to address this issue, but many existing techniques are model-specific and designed for classification tasks, making them less effective for object detection and difficult for non-specialists to interpret. In this work we focus on model-agnostic XAI methods for object detection models and propose D-MFPP, an extension of the Morphological Fragmental Perturbation Pyramid (MFPP), which uses segmentation-based mask generation. Additionally, we introduce D-Deletion, a novel metric combining faithfulness and localization, adapted specifically to meet the unique demands of object detectors. We evaluate these methods on real-world industrial and robotic datasets, examining the influence of parameters such as the number of masks, model size, and image resolution on the quality of explanations. Our experiments use single-stage object detection models applied to two safety-critical robotic environments: i) a shared human-robot workspace where safety is of paramount importance, and ii) an assembly area of battery kits, where safety is critical due to the potential for damage among high-risk components. Our findings evince that D-Deletion effectively gauges the performance of explanations when multiple elements of the same class appear in the same scene, while D-MFPP provides a promising alternative to D-RISE when fewer masks are used.
keywords:
Explainable Artificial Intelligence , Safe Artificial Intelligence , Trustworthy Artificial Intelligence , Object Detection , Single-stage object detection, Industrial Robotics
[…]
[tecnalia]organization=TECNALIA, Basque Research and Technology Alliance (BRTA),addressline=Mikeletegi Pasealekua 2, city=Donostia-San Sebastian, postcode=20009, country=Spain
[…]
[deusto] organization=University of Deusto, postcode=20012, city=Donostia-San Sebastián, country=Spain
[…]
[upv]organization=University of the Basque Country (UPV/EHU),city=Bilbao, postcode=48013, country=Spain
1 Introduction
In recent years, Artificial Intelligence (AI) has emerged as a transformative force across various domains, especially in human-machine interaction, where it has enabled significant advancements in data-driven decision-making processes. Among these advances, object detection has become a key component, finding application in critical areas such as autonomous driving, security surveillance, industrial automation, and robotics Zou et al. (2023); Muhammad et al. (2020). State-of-the-art object detection models, including Faster-RCNN Ren et al. (2017), DETR Carion et al. (2020), and the YOLO series Terven et al. (2023), have demonstrated impressive performance in identifying and localizing objects within images. Despite their success, the adoption of these models in highly sensitive environments remains limited, particularly in domains where errors could result in serious consequences such as injury, equipment damage, or operational failures. One of the primary reasons for this hesitancy is the black-box nature of object detectors implemented as Deep Learning models, which to date amount to the majority of proposals in the literature. The internal activations of these are not inherently interpretable, making it challenging for end-users to trust the predictions issued by object detectors, especially in high-risk environments operating in open-world environments such as autonomous vehicles and industrial robotics.
In this context, the field of Explainable AI (XAI) Barredo Arrieta et al. (2020) aims to enhance the interpretability of AI systems by their audience and ultimately, to enhance the user’s trust in the output of AI-based systems. Leaving aside the category of transparent AI models (which are inherently interpretable and do not require any explanations for a user to understand how they work), explainability methods in XAI can be broadly categorized into white-box and black-box approaches. White-box XAI methods require access to the internal workings of the model, such as weights, activations, or gradients (e.g., Grad-CAM Selvaraju et al. (2017)). While these methods can provide powerful insights, they are often limited by their dependence on specific model architectures, making them difficult to generalize across different models and less accessible to users unfamiliar with AI research/tools. In contrast, black-box XAI methods treat the model as an opaque entity, providing explanations based solely on the model’s input-output behavior without requiring any access to its internal components. However, most black-box XAI methods are designed for classification tasks rather than for object detection Ribeiro et al. (2016); Lundberg and Lee (2017); Petsiuk (2018); Ali et al. (2023a).
While classification models produce a single label per image, object detection models must identify and localize multiple objects within an image. Therefore, they need to explain not only the class prediction for each detected object – what they detect– but also the spatial reasoning behind the bounding boxes that define the object’s location – where the object is positioned within the image. Balancing these dual aspects complicates the explanation process and requires more sophisticated techniques than those used for classification tasks.
In this paper, we address the gap in XAI methods for object detection by focusing on model-agnostic, black-box XAI techniques. We propose and evaluate novel black-box XAI methods and XAI metrics that are specifically tailored for object detection models, without requiring access to internal model details. Our proposed methods are generalizable to object detection frameworks beyond those utilized in our experiments. Specifically, the contributions of this work can be summarized as follows:
  • We formally define a quantitative evaluation metric, D-Deletion, which extends the existing Deletion metric Bach et al. (2015); Petsiuk (2018) proposed for classification tasks. This metric is adapted to handle the unique challenges of object detection, including localization (as seen in Figure 4), which is of utmost importance when multiple instances of the same object appear in the same scene.
  • By using the similarity score of D-RISE Petsiuk et al. (2021), we analyze multiple mask generation methods’ performance and introduce D-MFPP, an extension of MFPP Yang et al. (2020) originally developed for classification tasks. D-MFPP utilizes segmentation-based mask generation to improve explanations for object detection models.
  • We analyze the impact of key parameters, such as image dimensions and the model sizes within the YOLOv8 architecture utilized in our experiments, which can significantly influence the quality of the resulting explanations.
  • Last but not least, we facilitate the broader adoption of the developed techniques for object detection in real-world use cases by releasing the code publicly in a repository https://github.com/aklein1995?tab=repositories  [^1] To be published upon article acceptance..
The remainder of this paper is structured as follows: in Section 2, we first review literature related to XAI for object detection. In Section 3, we provide the necessary background on object detection and XAI to familiarize the reader with the key concepts used in the definitions of D-RISE and Deletion. Next, Section 4 presents the experimental setup, including datasets, object detection training configuration, employed XAI methods, and evaluation metrics. In this section we also introduce our proposed D-MFPP method and D-Deletion metric. We discuss our results in Section 5. Finally, Section 6 concludes the paper with a summary of our key findings and directions for future research.
2 Related Work
Before proceeding with the materials and novel methods introduced in this work, we first pause briefly at XAI methods, focusing on those used for object detection tasks and put to practice in industrial applications:
XAI methods
As stated in the introduction, XAI offers insights into the procedure followed by an AI-based system to elicit their outputs, enabling end-users to understand and eventually trust the decisions output by the AI-based system grounded on objective data Ali et al. (2023b). To date, the majority of XAI methods are designed for models learned to address classification tasks. For instance, CAM-based methods like GradCAM Selvaraju et al. (2017), GradCAM++Chattopadhay et al. (2018) and Integrated Gradients Sundararajan et al. (2017) quantify and attribute the pixel-wise importance of a given input according to the gradients with respect a target class. Moreover, making use of backpropagation, LRP Montavon et al. (2019) calculates the contribution that a neuron has with neurons in consecutive layers to get relevance scores. In contrast, perturbation-based techniques work by occluding certain parts of the input and analyzing its impact in the predictions. Within this type of techniques, LIME Ribeiro et al. (2016), approximates a NN with an interpretable model; SHAP Lundberg and Lee (2017) assigns importance values to each input feature based on Shapley values; RISE Petsiuk (2018) generates saliency maps by probing the model with randomly masked versions of the input image; and MFPP Yang et al. (2020) generates masks by dividing the input image into multi-scale superpixels. Nonetheless, none of them have been explicitly extended for object detection tasks –with the exception of RISE, which has been adapted for this purpose– although techniques like SHAP can also be utilized for regression problems.
XAI methods for object detection models
In recent times, a scarcity of XAI approaches has been proposed to support the interpretability of complex object detection models. SODEx Sejr et al. (2021) is a method capable of explaining any object detection algorithm using classification explainers, demonstrating how LIME can be integrated within YOLOv4, a variant of the YOLO family of single-stage object detectors. Similarly, D-RISE Petsiuk et al. (2021) extends RISE’s mask generation technique by introducing a new similarity score that assesses both the localization and classification aspects of object detection models. More recently, D-CLOSE Truong et al. (2024) enhances D-RISE by producing less noisy explanations. Along with other methodological improvements, D-CLOSE uses multiple levels of segmentation in the mask generation phase. Other approaches focusing on hierarchical masking have been proposed. Concretely, GSM-NH Yan et al. (2022) evaluates the saliency maps at multiple levels based on the information of previous less fine-grained saliency maps, whereas BODEM Moradi et al. (2024) further extends this idea but focuses on an extreme black-box scenario where only object coordinates are available.
XAI methods for industrial applications
Although XAI is increasingly important in industrial settings to ensure safety, reliability and compliance, the adoption of XAI for object detection methods in industrial use cases has been limited to date Le et al. (2023); Kotriwala et al. (2021); Chen (2023). The vast majority of the works focus either on image classification, like Chen and Lee (2020) that utilizes Grad-CAM to interpret vibration signal images in the classification of bearing faults; time-series data, e.g. Serradilla et al. (2020) that presents the implementation and explanations of a remaining life estimator model; or tabular data, as in Ryo (2022) where SHAP is used to interpret and study the influence of soil and climate features on crop recommendations. Regarding XAI and object detection for industrial applications, we can find a few exemplary studies that expose the shortage of real-world use cases currently noted in this technological crossroads. In Naddaf-Sh et al. (2022), various object detection models are evaluated for their effectiveness in detecting weld characteristics in radiography images, with an emphasis on explainability and deployment on edge devices to assist workers. In the same sense, Sahatova and Balabaeva (2022) provides a comprehensive review and analysis of various XAI techniques applied to object detection tasks in computerized tomography imaging for medical purposes. Finally, Kirchknopf et al. (2022) demonstrates how to integrate Grad-CAM into the YOLO architecture and performs experiments in both public and private datasets of vehicle front collision and rear-view cameras.
3 Background
We now proceed by elaborating on key concepts needed to properly understand the details of the proposed D-MFPP technique and the D-Deletion metric that lie at the core of this work. Concretely, we provide fundamentals for object detection models (Subsection 3.1) and attribution-based for object detection
3.1 Object Detectors
Object detectors are crucial components in computer vision tasks, capable of identifying and localizing objects within an image. They can be broadly categorized into one-stage and two-stage detectors.
One-stage detectors
They directly predict bounding boxes and class probabilities from input images in a single pass. These detectors treat object detection as a simple regression problem, straight from image pixels to bounding box coordinates and class probabilities. To this end, they produce a dense grid of bounding box proposals and class probabilities in one step. Specifically, YOLO divides the input image into a grid and predicts bounding boxes and class probabilities for each grid cell. One-stage detectors are usually adopted for real-time applications due to their high speed and efficiency, although their accuracy tends to be lower when being compared with two-stage detectors.
Two-stage detectors
These models, among which Faster R-CNN Ren et al. (2017) can be considered to be the most representative one, follow a more complex approach that divides the detection process into two stages. In the first stage, a Region Proposal Network (RPN) generates a set of candidate object proposals (bounding boxes) from the input image. In the second stage, these proposals are refined and classified into different object categories by a second network. This second stage typically involves a more complex network, such as a convolutional neural network (CNN), which performs classification and further refinement of the bounding box coordinates. Two-stage detectors are known for their high accuracy and robustness in detecting objects of varying sizes and shapes, but they are generally slower than one-stage detectors.
Most detector networks, including Faster R-CNN and YOLO, produce a large number of bounding box proposals which are subsequently refined using confidence thresholding and Non-Maximum Suppression (NMS) to produce a set of finally detected objects in the image. Each bounding box proposal $d_{i}$ can be defined as follows:
$\mathbf{d}_{i}=\left[\mathbf{L}_{i},O_{i},\mathbf{P}_{i}\right]=\left[(x^{i}_{1},y^{i}_{1},x^{i}_{2},y^{i}_{2}),O_{i},(p^{i}_{1},\ldots,p^{i}_{C})\right],$ (1)
where $\mathbf{L}_{i}$ defines the bounding box corners $(x^{i}_{1},y^{i}_{1})$ and $(x^{i}_{2},y^{i}_{2})$; $O_{i}\in[0,1]$ refers to the probability that bounding box $L_{i}$ contains an object of any class; and $\mathbf{P}_{i}$ is a vector of probabilities $(p^{i}_{1},\ldots,p^{i}_{C})$ representing the probability that region $\mathbf{L}_{i}$ belongs to each of $C$ classes. As it can be observed in this above definition, unlike traditional classifiers which assign a single class label to an entire image, object detectors must simultaneously handle both classification and localization. This dual task involves predicting not only the class of each object but also its precise location within the image. The complexity of this task increases the difficulty of making these models interpretable.
3.2 Explainable Artificial Intelligence (XAI)
Despite the great performance exhibited by object detectors in manifold applications, their adoption in risk-sensitive scenarios is often hindered by a lack of trust and transparency by the user making decisions based on the detections issued by these models. As introduced previously, research on XAI produce techniques and methods that make the behavior and predictions of AI models understandable to humans without sacrificing performance Gunning et al. (2019). To this end, multiple XAI techniques have been proposed, which can be classified into four broad categories Ali et al. (2023b):
  • Scoop-based techniques focus on the extent of the explanation, providing either local explanations for specific predictions or global explanations for the overall model behavior.
  • Complexity-based methods consider the complexity of the model, with simpler, interpretable models offering intrinsic interpretability and more complex models requiring post-hoc explanations.
  • Model-based approaches distinguish between XAI methods that are specific to particular types of models, and those that are model-agnostic, capable of being applied to any model disregarding the specifics of their internals.
  • Methodology-based techniques are categorized by their methodological approach, such as backpropagation-based methods that trace input influences, or perturbation-based methods that alter inputs to observe changes in the output of the model.
Given that object detectors are typically complex neural networks, they fall under the complexity-based category, thereby requiring post-hoc explainability methods to explain their decisions. Among the various methodology-based techniques, attribution methods are commonly used to estimate the relevance of each pixel in an image for the detection task. Attribution methods are particularly important for object detection, where both localization and classification need to be explained.
Traditional attribution methods have been primarily developed for image classifiers Abhishek and Kamath (2022), which produce a single categorical output, making them less suited for object detectors. Object detectors, unlike classifiers, generate multiple detection vectors that encode not only class probabilities, but also localization information and additional metrics, such as objectness scores (see Section 3.1). Furthermore, techniques like NMS and confidence threshold filtering, which are used to refine bounding box proposals, add complexities that require a deeper understanding of the model’s internal workings, complicating the use of certain XAI methods, such as gradient-based approaches. Therefore, we focus on model-agnostic black-box XAI approaches, which are designed to be architecture-independent, and do not depend at all on the specifics of the model under target.
Among black-box XAI methods, perturbation-based approaches are commonly used due to their simplicity and effectiveness in revealing which parts of the input are most influential for the model’s predictions. Perturbation-based techniques offer a direct way to assess how changes to the input image affect the model’s output. By systematically altering or masking parts of the input image (using masks to generate perturbed samples), these methods allow inferring the importance of different regions based on the model’s input-output behavior.
The typical pipeline for perturbation-based XAI methods can be divided into three stages: (1) Data Preparation, (2) Model Assessment, and (3) Importance Computation. In the Data Preparation stage, masks are generated and applied to the image to create perturbed samples. The Model Assessment stage involves passing these perturbed images through the model to observe the changes in output. Finally, in the Importance Computation stage, the importance of each pixel is calculated by comparing the model’s outputs for the original and perturbed images. While the Model Assessment stage remains consistent across methods, with each perturbed image passed through the model, the Importance Computation varies depending on the XAI approach used. This can range from simple techniques like retraining a model (e.g., LIME) to more complex approaches. Since the effectiveness of these methods largely depends on how the perturbed images are generated, three mask generation algorithms are next described (Figure 1):
  • Sliding Window: This method, which is similar to the Occlusion technique proposed in Zeiler and Fergus (2014), systematically moves a window of fixed size across the image and sets the region within the window to a constant value (e.g., zero) to occlude that part of the image. By iteratively sliding the window across the entire image, we can assess the impact of each occluded region on the model’s output. The method requires specifying the window size, which determines the area of the image being occluded at each step, and the stride, which sets how much the window moves between iterations.
  • RISE: Randomized Input Sampling for Explanation (RISE) Petsiuk (2018) involves sampling $N$ binary masks of size $h\times w$, which are smaller than the original image size $H\times W$. Each element in the mask is independently set to 1 with probability $p$ and to 0 with the remaining probability $1-p$. These masks are then upsampled to size $(h+1)\cdot C_{H}\times(w+1)\cdot C_{W}$ using bilinear interpolation, where $C_{H}\times C_{W}=\left\lfloor H/h\right\rfloor\times\left\lfloor W/w\right\rfloor$. The upsampled masks are cropped to the original image size $H\times W$ with uniformly random offsets ranging from $(0,0)$ to $(C_{H},C_{W})$. This method creates a diverse set of masks that cover different parts of the image, allowing for a comprehensive evaluation of the importance of various regions.
  • MFPP: The so-called Morphological Fragmental Perturbation Pyramid (MFPP) Yang et al. (2020) method divides the input image into multi-scale fragments and perturbs them randomly. In this sense, it is similar to RISE, but instead of perturbing elements of the generated masks with dimension $h\times w$, MFPP defines regions according to segmentations at different scales. Depending on the number of defined fragments, the regions would be more fine-grained yet more time-consuming. The segments are dependent on each image, requiring the creation of new masks for every image.
Figure 1: Example of three masks generated using Sliding Window (top), RISE (middle), and MFPP (bottom). MFPP masks are dependent on the image at the input of the model. In this case, we consider a sample from the battery assembly dataset detailed in Section 4.
4 Materials and Methods
This section describes the industrial robotics use cases providing the datasets (Subsection 4.1), XAI methods (Subsection 4.3) and the explanation quality metrics (Subsection 4.4) considered in our work. The novel XAI technique and quality metrics proposed in this manuscript are also described in detail in Subsection 4.3.
4.1 Industrial Robotics Datasets under Consideration
The datasets used in this manuscript have been collected during the course of the ULTIMATE project, https://ultimate-project.eu/, which features two distinct real robotics use cases Kozik et al. (2024). The first dataset, from PIAP https://piap.lukasiewicz.gov.pl/), involves a collaborative workspace where a human and a robotic arm work together. The second dataset, provided by Robotnik https://robotnik.eu/, focuses on a battery assembly area, where a robotic arm assembles components for a battery kit.
Dataset 1: Human-Robot Dataset
This dataset consists of 96 images captured from three different cameras, as exemplified in Figure 2, with 32 images taken from each camera. The dataset includes two object classes: human and gripper. Importantly, each image in this dataset contains only a single object of each class, meaning a maximum of one human and one gripper per image. To ensure a diverse and representative sample, we applied feature extraction using ResNet He et al. (2015) to obtain embeddings for the entire dataset. The dimensionality of these embeddings was reduced using Principal Component Analysis (PCA), followed by K-means clustering (with $k=8$ clusters). From each cluster, four images were randomly selected, resulting in a final subset. The data were split into three sets: 72 images for training (75%), 6 for validation (6.25%), and 18 for testing (18.75%). To maintain consistency, we applied the same partitioning to the data from each camera. This resulted in 24 images for training, 2 for validation, and 6 for testing from each camera.
Figure 2: Dataset 1 (Human-Robot collaboration): Data are captured from cameras located in 3 different positions. All the images belonging to this dataset contain the faces blur to preserver anonymity.
Dataset 2: Battery Assembly Dataset
This dataset consists of 7 images, all captured from a bird’s-eye (top-down) view, showing a robotic arm assembling a battery kit, as shown in Figure 3. The dataset includes five distinct object types: individual battery, bms_a, bms_b, battery holder, and unknown object. In contrast to the Human-Robot Dataset, each image in the Battery Assembly Dataset may contain multiple objects of the same class, such as several individual batteries in a single scene.
Refer to caption    Refer to caption
Figure 3: Dataset 2 (Battery Assembly kit): The setup where a robotic arm would assemble the kit based a bird-eye view of the table where all component are expected to be; (left) a theoretical setup; (right) an actual sample.
It is worth noting that XAI techniques can be applied to any type of data. When applied to training data, they help reveal what the model has learned to focus on during training. When applied to test data, they provide insight into how well the model generalizes to new, unseen examples. For the Human-Robot Dataset, XAI explanations were applied exclusively to the test images, allowing us to assess the model’s behavior on unseen data. However, for the Battery Assembly Dataset, given the limited number of images (only 7), XAI explanations were applied to the entire dataset.
4.2 Object Detection Model: YOLOv8
Among the possible object detector models, we selected one of the state-of-the-art options, YOLOv8, due to its numerous advancements over previous versions and its robust performance in object detection tasks Terven et al. (2023). YOLOv8 Reis et al. (2024) integrates a novel combination of Feature Pyramid Network (FPN) and Path Aggregation Network (PAN) architectures, enhancing its ability to detect objects at various scales and resolutions. The FPN gradually reduces the spatial resolution of the input image while increasing feature channels, facilitating multi-scale object detection. The PAN architecture further aggregates features from different levels through skip connections, improving the detection of objects with diverse sizes and shapes. Additionally, YOLOv8 introduces an anchor-free detection mechanism that directly predicts the center of an object (instead of the offset from a known anchor box), reducing the number of box proposals and speeding-up the post-processing. Furthermore, it was trained with larger and more diverse datasets including the popular COCO dataset, improving its performance across a wider range of images.
YOLOv8 was developed and released by Ultralytics, and although the model and its weights are open-source, most users are expected to utilize the Ultralytics framework for its enhanced usability. However, unlike previous YOLO releases where the probability for each class per predicted box was accessible, in YOLOv8, the Ultralytics API outputs only the probability for the class with the highest confidence in each box [^2] https://github.com/ultralytics/ultralytics/issues/2863
https://github.com/ultralytics/ultralytics/issues/4908
. Consequently, by default, YOLOv8 outputs:
$\mathbf{d}_{i}=\left[\mathbf{L}_{i},O_{i},C_{i}\right]=\left[(x^{i}_{1},y^{i}_{1},x^{i}_{2},y^{i}_{2}),O_{i},C_{i}\right],$ (2)
where $\mathbf{L}_{i}=(x_{1}^{i},y_{1}^{i},x_{2}^{i},y_{2}^{i})$ represents the coordinates of the bounding box, $O_{i}$ denotes the objectness score, and $C_{i}$ corresponds to the predicted class label for the object within the bounding box, which differs with respect to the outputs shown in Expression (1).
4.3 Explainability Methods
We evaluate four popular methods for generating visual explanations of black-box models: LIME, RISE, D-RISE, and D-MFPP. The first two methods, LIME and RISE, were originally developed for image classifiers but can be adapted to object detectors. However, they primarily focus on explaining classification aspects and are not capable of addressing localization characteristics. In contrast, D-RISE is one of the first XAI methods specifically designed for object detectors, providing explanations that encompass both classification and localization. Additionally, we extend the existing MFPP method (originally tailored for classifiers) into a version suitable for object detection, which we refer to as D-MFPP. In what follows we briefly describe them, flowing into a description of the proposed D-MFPP approach:
  • LIME was originally designed to explain the predictions of any classifier by approximating it locally with an interpretable model. To explain the prediction for an input image $I$, LIME fits an interpretable model $g$ (e.g., a linear model) to approximate the behavior of the black-box model $f$ locally around $I$. The similarity between the original image and the perturbed samples is measured using a kernel function $\pi_{I}(z)$. When image explanations are targeted, LIME groups contiguous pixels into superpixels based on similar features they represent. This approach allows LIME to measure the importance of regions in the image rather than individual pixels, making the explanations more interpretable.
  • As introduced in the previous section, RISE Petsiuk (2018) was originally designed for deep neural networks that take images as input and output a class probability (e.g., a classifier like ResNet-50). It generates saliency maps that indicate the importance of each pixel by applying randomly generated binary masks $M_{i}$ to the input image $I$ and observing the changes in the model’s output $f(I\odot M_{i})$. In RISE, $N$ binary masks $M_{i}\in\{0,1\}^{h\times w}$ are generated (as explained in Section 3.2). These masks are then applied to the input image $I$ to generate masked images $I_{i}^{\prime}=I\odot M_{i}$, where $\odot$ denotes element-wise multiplication. The model is evaluated on each masked image $I_{i}^{\prime}$ to obtain the outputs $f(I\odot M_{i})$. The importance score for each pixel $(x,y)$ is then calculated as the weighted sum of the outputs:
    $S_{I,f}(x,y)=\frac{1}{N}\sum_{i=1}^{N}f(I\odot M_{i})\cdot M_{i}(x,y)$ (3)
    where the weights $M_{i}(x,y)$ represent the value of mask $i$ at pixel $(x,y)$. The intuition behind RISE is that $f(I\odot M_{i})$ would be high when pixels preserved by mask $M_{i}$ are important. Although this is true when having infinite diverse masks, in practice RISE calculates each pixel’s importance empirically by Monte Carlo sampling. Therefore, RISE largely depends on the number of masks ($N$) and how they are generated (i.e., is sensitive to the selected probability $p$ and resolution $s$).
4.3.1 D-RISE and Proposed D-MFPP Approach
Unlike the other two approaches originally designed for classifiers that measure solely classification aspects, D-RISE (Detector Randomized Input Sampling for Explanation) Petsiuk et al. (2021) was designed to explain both the classification and localization of a detection. In this sense, D-RISE extends RISE by producing saliency maps specifically for object detectors. As previously seen in Section 3.1, the output given by an object detector differs from the probability vector given by a classifier, obtaining localization information $L_{i}$, an objectness score $O_{i}$ and the probability of classifying each bounding box to any of the considered classes $P_{i}$. As a consequence, Expression (3) used by RISE is replaced in D-RISE with a new similarity score, given by:
$S_{I,f}(\mathbf{d}_{t},\mathbf{d}_{j})=s_{L}(\mathbf{d}_{t},\mathbf{d}_{j})\cdot s_{P}(\mathbf{d}_{t},\mathbf{d}_{j})\cdot s_{O}(\mathbf{d}_{t},\mathbf{d}_{j}),$ (4)
where $s_{L}=IoU(\mathbf{L}_{t},\mathbf{L}_{j})$, $s_{P}=\mathbf{P}_{t}\cdot\mathbf{P}_{j}/(||\mathbf{P}_{t}||\cdot||\mathbf{P}_{j}||)$, and $s_{O}=O_{j}$. In this formulation, $s_{L}$ represents the spatial proximity of the bounding boxes encoded by the target detection $\mathbf{d}_{t}$ and the proposal $\mathbf{d}_{j}$, measured using the Intersection over Union (IoU); the term $s_{P}$ evaluates the similarity between the class probabilities of the target detection and the proposal using cosine similarity; and $s_{O}$ incorporates the objectness score of the proposal $O_{j}$. It is important to note that for a detection target $\mathbf{d}_{t}$ there would potentially be more than one detection proposals $\mathbf{d}_{j}$. Therefore, we would have multiple $S_{I,f}(\mathbf{d}_{t},\mathbf{d}_{j})$. As explained in D-RISE, the explanations consider only the detection with maximal score for each mask:
$S_{I,f}\left(\mathbf{d}_{t},f(M_{i}\odot I)\right)=\max_{\mathbf{d}_{j}\in f(M_{i}\odot I)}S_{I,f}(\mathbf{d}_{t},\mathbf{d}_{j}).$ (5)
Given the YOLOv8 outputs explained in Section 4.2, which do not provide the class probability vector $P_{i}$ without modifying its architecture (an approach we want to avoid within the scope of this paper), we must adapt the similarity score to only consider $s_{L}$ and $s_{O}$. Consequently, the modified similarity score can be expressed as:
$S_{I,f}(\mathbf{d}_{t},\mathbf{d}_{j})=s_{L}(\mathbf{d}_{t},\mathbf{d}_{j})\cdot s_{O}(\mathbf{d}_{t},\mathbf{d}_{j})=IoU(\mathbf{L}_{t},\mathbf{L}_{j})\cdot O_{j}.$ (6)
This adjustment allows still utilizing D-RISE effectively for generating saliency maps with the default YOLOv8 model, focusing on the spatial and objectness aspects of detections, while maintaining the integrity of the model’s original architecture.
Similarly, we can adopt this similarity score but apply it with a different mask generation process. The MFPP method introduced in Section 3.2, originally designed for classification tasks, can be extended by applying Equation (6), resulting in D-MFPP. To the best of our knowledge, no previous work has proposed this variant of MFPP for object detection tasks.
4.4 Metrics
Evaluating the performance of attribution-based explainability methods for image data involves assessing how well the generated relevance heatmaps highlight important regions of the input image that contribute to the model’s decision. Generally, according to Hedström et al. (2023), explanation quality metrics can be grouped into six categories based on their logical similarity: faithfulness, robustness, localization, complexity, randomization, and axiomatic metrics. In this study, we focus on two of these categories that are particularly relevant to object detection: localization (Subsection 4.4.1) and faithfulness (Subsection 4.4.2).
4.4.1 Localization
Localization metrics evaluate whether the explainable evidence is centered around a region of interest (RoI) defined by a bounding box, segmentation mask, or a cell within a grid. These metrics aim to verify if the saliency maps correctly highlight the areas in the image that contain the object of interest. Among them, our experiments will consider:
  • Pointing Game (PG), which is a human evaluation metric introduced in Zhang et al. (2018). If the highest saliency point lies inside the human-annotated bounding box of an object, it is counted as a hit. The PG accuracy is given by:
    $\text{PG}=\frac{\#Hits}{\#Hits+\#Misses},$ (7)
    which is averaged over all categories in the dataset.
  • Energy-based Pointing Game (EBPG) Wang et al. (2020), which measures the proportion of activations within the given bounding box relative to the whole activation in the image. It assesses how much of the model’s activation energy is concentrated within the predefined region of interest. Formally:
    $\text{EBPG}=\frac{\sum_{(x,y)\in\text{bbox}}S_{I,f}(x,y)}{\sum_{(x,y)\in\text{bbox}}S_{I,f}(x,y)+\sum_{(x,y)\notin\text{bbox}}S_{I,f}(x,y)},$ (8)
    where $S_{I,f}(x,y)$ represent the saliency score at pixel $(x,y)$, $\sum_{(x,y)\in\text{bbox}}S_{I,f}(x,y)$ represents the sum of activation values within the bounding box, and $\sum_{(x,y)\notin\text{bbox}}S_{I,f}(x,y)$ represents the sum of activation values outside the bounding box.
4.4.2 Faithfulness
Figure 4: Illustration of a collaborative workspace featuring two humans and a robotic arm. The first row shows the original image. The second row displays the image with the 10% most important pixels removed for each human, as identified by an XAI method. In the third row, the Deletion metric curve, which only considers class type, shows a high probability score even when the primary human is largely occluded by the other person. The fourth row presents the D-Deletion metric curve, which incorporates a localization component, providing a more accurate measure of explanation importance by considering the positions of entities within the image. A lower area under the curve indicates a better explanation.
Metrics accounting for faithfulness quantify to what extent explanations follow the predictive behavior of the model, asserting that more important features play a larger role in model outcomes. These metrics focus on understanding the causal relationship between input features and the model’s output by systematically altering the features and observing the changes in predictions. Among them:
  • Deletion: Inspired by the work by Bach et al. (2015), the Deletion metric was proposed in RISE Petsiuk (2018). This metric measures a decrease in the probability of the predicted class as more and more important pixels are removed, where the importance is obtained from the saliency map. A sharp drop, and thus a low Area Under the probability Curve (AUC, as a function of the fraction of removed pixels), indicates a good explanation. Given the importance score for each pixel calculated by any XAI method, $S_{I,f}$, we can formulate the Deletion metric as:
[…]
    where $I$ is the original image, $M_{k}$ represent a mask with the the $k$-th most important pixels removed sorted by $S_{I,f}$, $Pr(f\left(I\odot M_{k}\right)=c)$ represents the probability of model $f$ predicting that the bounding box belongs to class $c$, and AUC($\cdot$) computes the area under the curve for the $K$ predictions.
  • Minimum Subset: It follows the same logic as Deletion, but instead of determining the AUC, it considers the required number of pixels that make the prediction to change Gevaert et al. (2024). Given the importance score for each pixel ($S_{I,f}$), Min-Subset is defined as the smallest subset of pixels that needs to be removed to change the model’s prediction. Mathematically,
[…]
    $\min\left\{k\in\{1,2\ldots,K\}:\>f\left(I\odot M_{k}\right)\neq f\left(I\right)\right\},$ (10)
    where $f(I\odot M_{k})$ represents the class label assigned by the model $f$ after passing the image $I$ with the top $k$ most important pixels removed, and $f(I)$ is the class label predicted for the original image.
4.4.3 Proposed D-Deletion and D-Minimal Subset Metrics
Algorithm 1 Deletion Metric’s Pseudocode for Object Detector
0:  Image $I$, saliency map $S_{I,f}$, number of steps $K$, target detection $\mathbf{d}_{t}$, target class $C_{t}$
1:  Initialize $S\leftarrow[]$
2:  for $k=1$ to $K$ do
3:     $M_{k}\leftarrow S_{I,f}$ removing the top $k$ most important pixels
4:     Apply mask $M_{k}$ to image $I$
5:     Forward pass through the model $f$ and obtain the bounding box proposals $\mathbf{d}_{j}=[\mathbf{L}_{j},O_{j},C_{j}]=f(I\odot M_{k})$
6:     Initialize list of proposals: proposals [$\leftarrow$[]← [ ]
7:     for each bounding box $\mathbf{d}_{j}$ predicted by the model $f$ do
8:        if $C_{j}=C_{t}$ then
9:           proposals.append($O_{j}$) % For D-Deletion 10:           if  $IoU(\mathbf{d}_{t},\mathbf{d}_{j})>\gamma$  then 11:              proposals.append($O_{j}$) 12:           end if
13:        else
14:           proposals.append(0)
15:        end if
16:     end for
17:     Insert the maximum score within the deletion buffer: $S\leftarrow S\cup\max(\texttt{proposals})$
18:  end for
19:  Calculate the Deletion metric as the area under the curve: $D=\text{AUC}(S)$
20:  return  Deletion score $D$
Originally, Deletion was designed for classifiers. However, with object detectors, multiple detections in a single image can occur. Although D-RISE stated the necessity to adapt this metric for object detectors Petsiuk et al. (2021), no formal definition can be found in the literature. Therefore, considering the importance of this issue in real use cases, we formally re-define Equation (9) in two manners:
  1. 1.
    Deletion. Measures the explanation given the target class label $C_{t}$ (regardless if there is more than one element for a class) and iteratively removes the top $k$ most important pixels:
    $\text{Deletion}(I,S,C_{t})=$
    $AUC\left(\left\{\max_{\mathbf{d}_{j}^{k}}\left[O_{j}^{k}\cdot\mathds{I}\{C_{j}^{k}=C_{t}\}\right]\right\}^{K}_{k=1}\right),$ (11)
    The model $f(\cdot)$ takes as input the masked image $I\odot M_{k}$ and outputs a set of bounding box proposals $\mathbf{d}_{j}^{k}=[\mathbf{L}_{j}^{k},O_{j}^{k},C_{j}^{k}]$. The indicator function $\mathds{I}\{C_{j}^{k}=C_{t}\}$ equals 1 if the predicted class $C_{j}^{k}$ matches the target class $C_{t}$, and 0 otherwise. The term $\max_{\mathbf{d}_{j}^{k}}\left[O_{j}^{k}\cdot\mathds{I}\{C_{j}^{k}=C_{t}\}\right]$ selects the maximum objectness score $O_{j}^{k}$ for the bounding boxes where the predicted class matches the target class. The AUC is then computed over the set of prediction scores for the $K$ steps, where at each step the most important pixels are progressively removed.
  2. 2.
    D-Deletion. While the standard Deletion metric evaluates the impact of pixel removal on a class prediction, it lacks the ability to account for spatial localization, which is essential in object detection tasks where multiple instances of the same class can appear. D-Deletion addresses this limitation by focusing on a specific target bounding box $\mathbf{d}_{t}$, considering both the class information, $C_{t}$, and IoU between the target and other detected proposals, $\mathbf{d}_{j}^{k}$. This ensures that the metric not only measures faithfulness but also takes localization into account, providing more precise explanations in situations where different objects of the same class coexist. Mathematically is expressed as:
[…]
    $AUC\left(\left\{\max_{\mathbf{d}_{j}^{k}}\left[O_{j}^{k}\cdot\mathds{I}\{C_{j}^{k}=C_{t}\}\cdot\mathds{I}\{IoU(\mathbf{d}_{t},\mathbf{d}_{j}^{k})>\gamma\}\right]\right\}^{K}_{k=1}\right)$ (12)
    where $\gamma$ is a threshold. As a consequence, when multiple elements of the same class are in an image, D-Deletion will only consider those proposals $\mathbf{d}_{j}^{k}$ predicted by the model that have a predefined IoU with the target bounding box $\mathbf{d}_{t}$.
The difference between Deletion and D-Deletion is illustrated in Figure 4. This figure highlights how D-Deletion distinguishes between different objects of the same class by incorporating localization information, leading to more refined and accurate explanations ($\downarrow$ AUC in the Figure’s last row) when multiple objects of the same class are detected in an image. For the sake of clarity, we provide the pseudocode of Deletion in Algorithm 1, where the main difference with respect to D-Deletion are lines 9 to 11.
Lastly, akin to how Min-Subset is related to Deletion, D-Min-Subset is associated with D-Deletion. Consequently, D-Min-Subset considers both the class type and the IoU to determine the number of pixels required to make the prediction to change:
[…]
$\min\left\{k\in\{1,2\ldots,K\}:\>C_{j}^{k}\neq C_{t}\text{ or }IoU(\mathbf{d}_{t},\mathbf{d}_{j}^{k})<\gamma\right\},$ (13)
where $C_{j}^{k}$ represents the predicted class label for detection $j$ when passing the masked image $I\odot M_{k}$ through the model $f$, with $\mathbf{d}_{j}^{k}=f(I\odot M_{k})$ being the set of detections after removing the top $k$ most important pixels. In this context, D-Min-Subset depends on two conditions: (1) the class probability labels $C_{j}^{k}$ for the predicted bounding box $\mathbf{d}_{j}^{k}$ must no longer match the target class $C_{t}$, or (2) the IoU between the target bounding box $\mathbf{d}_{t}$ and the predicted bounding box $\mathbf{d}_{j}^{k}$ falls below the threshold $\gamma$. The minimum $k$ is identified as the step where either of these conditions is first met.
5 Experiments and Results
Contrarily to most studies in the XAI literature that primarily focus on benchmark datasets, our research work focuses on assessing the explainability of object detectors in real-world industrial data. In this context, to evaluate the effectiveness of explanations, we formulate four key research questions to answer them with empirical evidence:
  • RQ1. Which XAI method provides the most reliable and insightful explanations for object detection models?
  • RQ2. Does the D-Deletion metric enhance the trustworthiness of XAI outputs when multiple objects of the same class are present in the image?
  • RQ3. How does the mask generation process influence the quality of explanations, particularly when using similarity scores for object detection? How does D-MFPP behave?
  • RQ4. Do different image dimensions impact the explanations generated by XAI methods? Do models of varying sizes (large, medium, small, nano) focus on different regions of the image in their explanations?
Next, we outline the hyperparameters used across our experiments to ensure consistency in training and evaluation. For both datasets, models were trained using the YOLOv8 architecture for a total of 100 epochs. The image size (imgsize) was set to the largest dimension of the input image (e.g., $1280720\times 1280\xrightarrow{}$ 1280), and data augmentation techniques such as random horizontal flipping and color jitter were applied. For consistency, the default Ultralytics settings were used wherever applicable. In the case of LIME, we use the baseline implementation of Ribeiro et al. (2016), where we adopt the SLIC segmentation algorithm Achanta et al. (2010) (with 100 segments) and generated 1000 samples to assess the quality of the produced explanations. For RISE and D-RISE, we employed 5000 masks with a probability of 0.25 and a resolution of $16\times 16$ to produce the saliency maps. Lastly, for all object detection predictions, a confidence threshold of 0.7 was set to determine the validity of each detection.
In what follows we present and discuss on the results obtained to answer each of the RQ formulated above:
RQ1. Comparison between XAI methods
Figure 5: Heatmaps obtained by applying RISE (left) and D-RISE (right) for the detection of a human in the Human-Robot Dataset.
Figure 6: Heatmaps generated in a scene of the Battery Assembly dataset for two target classes: individual battery (top row) and battery holder (bottom row). The first and second columns shows the saliency maps obtained using LIME and RISE, independent of the number of objects of the same class in the image. Columns 3,4 and 5 display heatmaps generated using D-RISE for three different individual elements of the same class.
In the Human-Robot dataset, the comparison between LIME, RISE, and D-RISE, as shown in Table 1, reveals distinct strengths across different metrics (Section 4.4). LIME performs good in terms of localization, with higher PG and EBPG scores (100% and 18.60%, respectively) compared to D-RISE (93.75% and 10.74%). This indicates that LIME generates more localized saliency maps, focusing closely on the bounding boxes of detected objects. However, this superior performance is partly due to the size of the object being analyzed. LIME’s superpixel generation is better suited for larger objects (e.g., human), as larger regions of the image can be grouped effectively into meaningful segments, leading to higher localization scores. This advantage also applies to classification, where larger objects allow LIME to better preserve relevant features for detection. Conversely, for smaller objects (e.g., gripper), LIME struggles when compared to the other methods, as reflected by its worse performance metrics in those cases.
In contrast, RISE and D-RISE are less sensitive to object size, making them more robust across different object scales, which is evident in their better performance on smaller objects like the gripper. They achieve Deletion scores of $0.2636$ and $0.2777$, respectively, compared to LIME’s $0.4688$. When considering the overall performance across classes, RISE, with a Deletion score of $0.2232$ and D-Deletion of $0.0651$, shows improvement over LIME in classification-related tasks but still lags behind D-RISE, which achieves the lowest Deletion ($0.2016$) and D-Deletion ($0.0438$) scores. Although D-RISE offers the best balance between classification and localization, the difference between RISE and D-RISE is minimal in this dataset, where each image contains only a single object per class. As a result, as shown in Figure 5, their heatmaps are very similar to each other, both highlighting the human head. However, D-RISE eliminates less relevant areas more effectively.
Table 1: Quantitative metrics of LIME, RISE and D-RISE for the Human-Robot dataset. The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each class and the overall average. $\downarrow$ indicates that lower values are better, whereas $\uparrow$ indicate that higher values are better. The best results for each metric and object type are highlighted in gray, while the best average values are written in bold.
XAI Method Object Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
LIME Human 0.0759 0.0632 4.2703 4.2703 1.0000 34.984
Gripper 0.4688 0.0324 1.3859 1.3859 1.0000 2.2270
Average 0.2723 0.0478 2.8281 2.8281 1.0000 18.6060
RISE Human 0.1827 0.1241 9.0108 9.0108 0.7500 19.0824
Gripper 0.2637 0.0060 0.6355 0.63 1.0000 1.0542
Average 0.2232 0.0651 4.8232 4.8232 0.875 10.0683
D-RISE Human 0.1255 0.0818 5.7335 5.7335 0.8750 20.4061
Gripper 0.2777 0.0059 0.6091 0.6091 1.0000 1.0815
Average 0.2016 0.0438 3.1713 3.1713 0.9375 10.7438
In the results obtained over the Battery Assembly dataset (Table 2), a similar pattern can be noticed. LIME excels at localization with an average EBPG of 16.03%, while RISE and D-RISE perform better in retaining key classification features. Since this dataset includes multiple objects of the same class (e.g., multiple batteries), both LIME and RISE, which are not designed to handle multiple detections of the same class, expose severe limitations. RISE, with a D-Deletion score of $0.1474$, preserves key features better than LIME, but is outperformed by D-RISE, which achieves a score of $0.0344$. D-RISE also shows the highest PG score (97%), performing significantly better than LIME (76.85%) and RISE (66.95%).
Table 2: Quantitative metrics of LIME, RISE and D-RISE for the Battery Assembly dataset. The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each object and the overall average. The best results for each metric and object type are highlighted in gray, while the best average values are marked in bold.
XAI Method Object Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
LIME indiv batt 0.7549 0.2806 44.3412 14.6756 0.3188 2.5347
bms a 0.0184 0.0181 0.8784 0.8784 1.0000 30.7573
bms b 0.0125 0.0125 0.8321 0.8321 1.0000 16.5440
unknown object 0.0624 0.0245 2.0342 2.0342 1.0000 20.1071
batt holder 0.1498 0.0849 6.8115 4.3766 0.5238 10.2087
Average 0.1996 0.0841 10.9795 4.5594 0.7685 16.0304
RISE indiv batt 0.7659 0.4359 81.2344 32.0482 0.0144 1.2558
bms a 0.0190 0.0190 1.9417 1.9417 1.0000 3.0409
bms b 0.0217 0.0217 2.1266 2.1266 1.0000 2.4561
unknown object 0.5333 0.2008 3.8372 3.8372 1.0000 5.5780
batt holder 0.0902 0.0595 7.9519 5.5632 0.3333 3.3681
Average 0.2860 0.1474 19.4184 9.1034 0.6695 3.1398
D-RISE indiv batt 0.6214 0.0311 35.7678 2.7725 1.0000 2.0546
bms a 0.0181 0.0181 1.9880 1.9880 1.0000 3.2955
bms b 0.0128 0.0116 1.3407 1.3407 1.0000 2.8876
unknown object 0.0879 0.0485 3.7448 4.2071 0.8571 7.6831
batt holder 0.4839 0.0626 21.7291 5.1009 1.0000 4.9635
Average 0.2448 0.0344 12.9141 3.0819 0.9714 4.1768
Overall, when dealing with datasets containing only one object per class, the differences between LIME, RISE, and D-RISE are relatively small in quantitative terms. However, when multiple objects of the same class appear in a given input image, D-RISE clearly dominates over the rest of techniques. As illustrated in Figure 6, D-RISE generates coherent heatmaps for each detected object in the Battery Assembly dataset, whereas LIME and RISE provide a global saliency map for the entire class. By combining the individual saliency maps from D-RISE, a more accurate and object-specific explanation can be produced. This also highlights the limitations of LIME and RISE when applied to multiple objects, as their global saliency maps do not differentiate between individual instances.
Class XAI $N=500$ $N=1,000$ $N=5,000$ $N=10,000$
Human
D-RISE
Refer to caption Refer to caption Refer to caption Refer to caption
D-MFPP
Refer to caption Refer to caption Refer to caption Refer to caption
Gripper
D-RISE
Refer to caption Refer to caption Refer to caption Refer to caption
D-MFPP
Refer to caption Refer to caption Refer to caption Refer to caption
indiv batt
D-RISE
Refer to caption Refer to caption Refer to caption Refer to caption
D-MFPP
Refer to caption Refer to caption Refer to caption Refer to caption
bms a
D-RISE
Refer to caption Refer to caption Refer to caption Refer to caption
D-MFPP
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Heatmaps obtained with D-RISE and D-MFPP using varying numbers of masks (500, 1000, 5000, 10000). Saliency maps are presented for two scenes, one from each dataset, demonstrating the evolution of heatmaps for a single element as the number of masks increases, highlighting the effect on explanation quality and focus.
RQ2. D-Deletion metric for scenes with multiple objects of the same class
As a secondary observation in the experiments of RQ1, the D-Deletion metric is specifically designed to overcome the limitations of traditional deletion metrics, particularly when multiple objects of the same class are present in an image.
In the Battery Assembly dataset, where several instances of the same class (e.g., indiv batt) appear, D-Deletion demonstrates clear advantages. By inspecting Table 2, RISE, while performing reasonably well with an average Deletion score of $0.2860$, it still obtains a relatively high D-Deletion score of 0.1474, suggesting that it struggles to differentiate between the contributions of individual objects. In contrast, D-RISE, which obtains an average Deletion score of 0.2448, outperforms RISE with a D-Deletion score of 0.03444. This highlights D-RISE’s ability to isolate and preserve key features for each object, providing more trustworthy, object-specific explanations rather than broad, class-level insights.
The Min-Subset and D-Min-Subset metrics, which measure the minimal proportion of pixels needed to disrupt a detection, reinforce these findings. In the Human-Robot dataset, Table 1, where only one object per class appears, the differences between Deletion and D-Deletion scores are minor, and the Min- Subset and D-Min-Subset values are close to each other. However, in the Battery Assembly dataset, where the differences between Deletion and D-Deletion are more substantial and multiple objects of the same class co-occur in the same image, the Min-Subset ($12.9141$) and D-Min-Subset ($3.0819$) values also diverge significantly.
RQ3. Influence of the Mask Generation Strategy
Figure 8: Explanations of a scene using different stride and window size configurations when using D-Sliding Window (combination of mask generation explained in Section 3.2 and Equation (6)). (Left) Stride of 16; (Right) Stride of 8; (Top) Window size of 32; (Bottom) Window size of 64.
When comparing XAI approaches for object detection tasks configured with different mask generation techniques, the results in Tables 3 and 4 initially suggest that D-Sliding Window performs the best in almost all metrics. However, as noted in the captions, this is only in cases where explanations were provided. For the Human-Robot dataset (Table 3), regardless of the window size and stride, D-Sliding Window failed to provide explanations for larger objects, such as humans, and only provided meaningful explanations for smaller objects like gripper instances. Similarly, in the Battery Assembly dataset (Table 4), D-Sliding Window struggled with large objects when using a smaller window size (w=32), which led to higher scores in classification metrics. Even with an increased window size (w=64), some objects were still not detected, making it difficult to fairly compare with D-RISE and D-MFPP, both of which provided explanations for all detection proposals.
In contrast, D-RISE consistently performs best across all detection proposals when explanations are provided, particularly in terms of the D-Deletion metric, with D-MFPP as a close second. D-MFPP manages to better distinguish important regions by de-emphasizing less relevant areas, as reflected in the PG and EBPG localization scores. Yet, due to the superpixel segmentation approach used by D-MFPP, some relevant features are grouped with less important ones (see Figure 7), causing a slight drop in classification metrics. As a result, D-MFPP trades off classification precision for improved localization and focus. D-Sliding Window, as previously noted, performs well on smaller objects but struggles to generate reliable explanations for larger objects. This limitation can be mitigated by using larger window sizes, which occlude larger portions of the image, thereby increasing the likelihood of capturing explanations for bigger objects, while smaller strides help produce more finely detailed explanations, as seen in Figure 8.
Influence of the number of masks
This secondary analysis within RQ3 confirms that the number of masks in use has a clear impact on the performance of the XAI method. Generally, more masks result in better outcomes. Interestingly, D-MFPP initially outperforms D-RISE in some metrics when using fewer masks (500 or 1,000), achieving a perfect PG score of 100% in the Human-Robot dataset and a D-Deletion score of 0.0505 with just 500 masks. However, as the number of masks increases, D-RISE slightly surpasses D-MFPP in classification-related metrics. This trend can be also noted in Figure 7, where (1) increasing the number of masks improves the quality of the saliency maps, and (2) D-MFPP provides cleaner maps with a lower number of masks.
Table 3: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Human-Robot dataset. Only the average score across all classes is presented. The best mask configuration for each XAI method is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
XAI Method Num Masks Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
D-Sliding Window (*) $w=32$, $s=8$ 0.2323 0.2323 24.1966 24.1966 1.0000 97.7128
$w=32$, $s=16$ 0.3009 0.3009 34.3990 34.3990 1.0000 96.5487
$w=64$, $s=8$ 0.2046 0.2002 20.7980 20.7980 1.0000 83.9141
$w=64$, $s=16$ 0.2039 0.1995 20.7362 20.7362 1.0000 83.5582
D-RISE 500 0.2645 0.0702 4.3520 4.3520 0.9062 10.7117
1000 0.2530 0.0519 4.2019 4.2019 0.9375 10.8382
5000 0.2016 0.0438 3.1713 3.1713 0.9375 10.7438
10000 0.2246 0.0389 3.2496 3.2496 0.9375 10.7094
D-MFPP 500 0.4137 0.0505 2.9608 2.9519 1.0000 15.9700
1000 0.4227 0.0678 5.8648 5.8648 1.0000 15.1166
5000 0.3944 0.0469 3.8146 3.8058 1.0000 16.0809
10000 0.3893 0.0471 3.6734 3.6734 1.0000 16.2579
*: D-Sliding Window scores have been computed excluding the cases where no explanations are produced for images with large-sized object instances.
Table 4: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Battery Assembly dataset. The best mask configuration for each XAI method is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
XAI Method Num Masks Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
D-Sliding Window (*) $w=32$, $s=8$ 0.5198 0.3258 53.9463 31.2219 0.9855 90.4736
$w=32$, $s=16$ 0.4700 0.3046 45.9925 24.7660 0.9788 84.5302
$w=64$, $s=8$ 0.2578 0.0245 26.4318 2.4062 0.9855 55.8151
$w=64$, $s=16$ 0.2570 0.0246 26.3537 2.4772 0.9884 54.8283
D-RISE 500 0.4307 0.0783 29.7843 4.9001 0.9971 3.8152
1000 0.3837 0.0701 26.4242 3.3443 0.9971 3.8154
5000 0.3645 0.0431 25.2766 3.1374 0.9971 3.8015
10000 0.3707 0.0342 23.3216 2.9789 0.9971 3.7883
D-MFPP 500 0.3327 0.0708 30.7555 5.6594 0.9942 6.2320
1000 0.3375 0.0630 30.0871 4.8843 0.9971 6.1938
5000 0.3042 0.0427 27.8461 3.4328 0.9971 6.1772
10000 0.3100 0.0409 29.6047 3.8185 0.9971 6.1680
*: D-Sliding Window scores have been computed excluding the cases where no explanations are produced for images with large-sized object instances.
RQ4. Image dimension variation and YOLOv8 complexity
Table 5: Explanations comparison for three image dimensions (720x1280, 360x640, and 180x320) and varying resolutions (16x16, 8x8, and 4x4) in the Human-Robot dataset.
Image dimensions Resolution Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
$720\times 1280$ $16\times 16$ 0.2016 0.0438 3.1713 3.1713 0.9375 10.7438
$8\times 8$ 0.0851 0.0512 4.0050 4.0050 1.0000 11.8188
$4\times 4$ 0.0541 0.0504 4.0684 4.0684 0.9166 12.4406
$360\times 640$ $16\times 16$ 0.4331 0.1697 10.6146 9.8644 0.7647 9.8634
$8\times 8$ 0.1889 0.1003 6.1750 6.1750 0.8823 11.0002
$4\times 4$ 0.1053 0.1013 8.2321 8.2321 0.9166 12.4447
$180\times 320$ $16\times 16$ 0.3668 0.0756 7.2829 7.2829 0.8957 10.5611
$8\times 8$ 0.2290 0.1206 10.1774 10.1774 1.000 11.6374
$4\times 4$ 0.1531 0.1232 11.6246 11.6246 0.9545 12.0230
Table 6: Impact of YOLOv8 model complexity (nano, small, medium, large) on the explanations for the same experimental setup.
Model size Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
Large 0.1735 0.1569 7.1964 7.1964 0.8333 10.0317
Medium 0.2016 0.0438 3.1713 3.1713 0.9375 10.7438
Small 0.3120 0.2294 5.3750 4.3060 0.7352 9.6248
Nano 0.1434 0.0930 4.1177 4.1177 0.9117 10.2061
We further analyze the importance of image dimensions in the quality of explanations. For this purpose, we interpolate the image dimensions and train a new object detector for those image dimensions. As reported in Table 5, larger image dimensions ($720\times 1280$) yield better results in terms of faithfulness metrics, degrading their value when decreasing the image resolution to $360\times 640$ and $180\times 320$. Nonetheless, PG and EBPG remain relatively stable across resolutions, suggesting that localization is less affected by image size than classification accuracy.
Additionally, Table 5 shows the effect of applying D-RISE with masks generated at varying resolutions. While no clear trend emerges in the classification metrics (with some instances of Deletion increasing while D-Deletion decreases), the localization metrics generally improve with lower mask resolutions. This result is expected, as lower-resolution masks produce less fine-grained explanations but tend to highlight a larger area around the target object. This pattern is clearly visualized in Figure 9.
Lastly, we investigate whether the use of different levels of parametric complexity of the YOLOv8 object detector affect the expected outcomes. When examining the results in Table 6, no clear insights can be drawn, as the best results were reported by the Medium and Nano models. Nonetheless, it is important to highlight the performance of the Nano model, which despite requiring half the inference time and an order of magnitude less memory for deployment, achieves competitive results with respect to the rest of counterparts in the table.
Class Resolution$=4\times 4$ Resolution$=16\times 16$
$720\times 1280$
Human
Refer to caption Refer to caption
Gripper
Refer to caption Refer to caption
$320\times 640$
Human
Refer to caption Refer to caption
Gripper
Refer to caption Refer to caption
$180\times 320$
Human
Refer to caption Refer to caption
Gripper
Refer to caption Refer to caption
Figure 9: Heatmaps generated with D-RISE ($p=0.25$ and resolutions of $4\times 4$ and $16\times 16$) on a scene from the Human-Robot scenario. The effect of varying input image dimensions can be observed.
6 Conclusion
This work has explored the performance of various XAI methods for object detection models, focusing on perturbation-based black-box explanation techniques like LIME, RISE, D-RISE applied to YOLOv8 in real-world object detection tasks in industrial setups. Our contribution is twofold: i) we have introduced D-MFPP, a method tailored for object detection that generates masks based on multi-level superpixels; and ii) we have proposed D-Deletion, a new quantitative metric that accounts for both class probability and localization when evaluating explanations.
Our results have demonstrated that D-RISE consistently outperforms LIME and RISE, particularly with the D-Deletion metric, which improves trustworthiness by delivering more focused, object-specific explanations. The mask generation process has been proven to play a key role in the quality of explanations, with more masks producing better explanations. The proposed D-MFPP technique has shown competitive performance with fewer masks, proving to be an efficient option in resource-constrained deployments. We have also found out that larger image dimensions improve classification metrics, while localization metrics remained stable across resolutions, suggesting that image size mainly affects classification. No clear patterns have emerged regarding model complexity; nevertheless, the nano YOLOv8 model, with its low computational and memory demands, showed promise for real-time industrial applications.
Limitations of the study
While this study provides valuable insights into the explainability of object detectors in real-world scenarios, several limitations must be acknowledged:
  • Limited data: The real-world datasets considered in the study, particularly the Battery Assembly use case, are limited in size (only 7 data instances). This may restrict the generalization of our findings. Future studies with larger datasets embracing the methodology herein followed could provide more robust conclusions.
  • Generalization of the object detector: Although the methods presented here are model-agnostic, the study primarily focuses on YOLOv8 (i.e., one stage detector). Extending this work to consider other object detector architectures, such as Faster R-CNN (i.e., two stage detector) or Detection Transformers (DETR, RT-DETR) could ease a broader understanding of XAI methods across different neural architectures for this modeling task.
  • Computation time: The computation time required to generate explanations can be significant, especially when using perturbation-based methods. Approaches that reduce the number of masks used, or optimize mask generation, could help mitigate this limitation and improve efficiency.
Future Research
We plan to explore hierarchical masking approaches Yan et al. (2022, 2024); Moradi et al. (2024) and feature fusion techniques Truong et al. (2024) that further refine pixel saliency across multiple levels, reducing masking efforts and enhancing the quality of the saliency maps. Moreover, we will explore techniques to reduce the computation time while maintaining performance and explanation quality Nguyen et al. (2024), which can be crucial in real-world critical applications wherein explanations have to be furnished fastly and presented to the human for its supervision and validation.
Acknowledgments
A. Andres, I. Laña and J. Del Ser receive support from the ULTIMATE project (ref. 101070162) funded by the European Commission under the HORIZON-CL4-DIS program (HORIZON-CL4-2021-HUMAN-01). J. Del Ser and I. Laña also acknowledge funding from the Basque Government (MATHMODE, IT1456-22).
Declaration of competing interest
The authors declare that they have no conflicts of interest regarding this work. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. The European Union can not be held responsible for them.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used OpenAI’s ChatGPT4 in order to enhance the clarity of the writing, improve the readability of the manuscript, and check for possible English language mistakes. After using this tool/service, the authors reviewed and edited the content as needed, and take full responsibility for the content of the publication.

### v2

@@
-On the Black-box Explainability of Object Detection Models for Safe and Trustworthy Industrial Applications
+On the Black-box Explainability of Object Detection Models for
 Abstract
@@
 However, their adoption in high-risk applications, where errors may cause severe consequences, remains limited.
-Explainable Artificial Intelligence (XAI) methods aim to address this issue, but many existing techniques are model-specific and designed for classification tasks, making them less effective for object detection and difficult for non-specialists to interpret.
-In this work we focus on model-agnostic XAI methods for object detection models and propose D-MFPP, an extension of the Morphological Fragmental Perturbation Pyramid (MFPP), which uses segmentation-based mask generation.
+Explainable Artificial Intelligence methods aim to address this issue, but many existing techniques are model-specific and designed for classification tasks, making them less effective for object detection and difficult for non-specialists to interpret.
+In this work we focus on model-agnostic explainability methods for object detection models and propose D-MFPP, an extension of the Morphological Fragmental Perturbation Pyramid (MFPP) technique based on segmentation-based masks to generate explanations.
 Additionally, we introduce D-Deletion, a novel metric combining faithfulness and localization, adapted specifically to meet the unique demands of object detectors.
@@
 Our experiments use single-stage object detection models applied to two safety-critical robotic environments: i) a shared human-robot workspace where safety is of paramount importance, and ii) an assembly area of battery kits, where safety is critical due to the potential for damage among high-risk components.
-Our findings evince that D-Deletion effectively gauges the performance of explanations when multiple elements of the same class appear in the same scene, while D-MFPP provides a promising alternative to D-RISE when fewer masks are used.
+Our findings evince that D-Deletion effectively gauges the performance of explanations when multiple elements of the same class appear in a scene, while D-MFPP provides a promising alternative to D-RISE when fewer masks are used.
 keywords:
 Explainable Artificial Intelligence , Safe Artificial Intelligence , Trustworthy Artificial Intelligence , Object Detection , Single-stage object detection, Industrial Robotics
+journal: Results in Engineering
 […]
@@
 (2016); Lundberg and Lee (2017); Petsiuk (2018); Ali et al.
-(2023a).
+(2023).
 While classification models produce a single label per image, object detection models must identify and localize multiple objects within an image.
@@
   • We analyze the impact of key parameters, such as image dimensions and the model sizes within the YOLOv8 architecture utilized in our experiments, which can significantly influence the quality of the resulting explanations.
-  • Last but not least, we facilitate the broader adoption of the developed techniques for object detection in real-world use cases by releasing the code publicly in a repository https://github.com/aklein1995?tab=repositories  [^1] To be published upon article acceptance..
+  • Last but not least, we facilitate the broader adoption of the developed techniques for object detection in real-world use cases by releasing the code publicly in a repository: https://github.com/aklein1995/drise_dmfpp_ddeletion.
 The remainder of this paper is structured as follows: in Section 2, we first review literature related to XAI for object detection.
@@
 As stated in the introduction, XAI offers insights into the procedure followed by an AI-based system to elicit their outputs, enabling end-users to understand and eventually trust the decisions output by the AI-based system grounded on objective data Ali et al.
-(2023b).
+(2023).
 To date, the majority of XAI methods are designed for models learned to address classification tasks.
@@
 We now proceed by elaborating on key concepts needed to properly understand the details of the proposed D-MFPP technique and the D-Deletion metric that lie at the core of this work.
-Concretely, we provide fundamentals for object detection models (Subsection 3.1) and attribution-based for object detection
+Concretely, we provide fundamentals for object detection models (Section 3.1) and XAI, with a focus on model-agnostic black-box methods to explain the predictions of object detection models (Section 3.2).
 3.1 Object Detectors
-Object detectors are crucial components in computer vision tasks, capable of identifying and localizing objects within an image. They can be broadly categorized into one-stage and two-stage detectors.
-One-stage detectors
+Object detectors are crucial components in computer vision tasks, capable of identifying and localizing objects within an image.
+They can be broadly categorized into single-stage and two-stage detectors.
+Single-stage detectors
 They directly predict bounding boxes and class probabilities from input images in a single pass.
-These detectors treat object detection as a simple regression problem, straight from image pixels to bounding box coordinates and class probabilities.
+Popular single-stage detectors, such as YOLO Terven et al.
+(2023), SSD Liu et al.
+(2016) and RetinaNet Ross and Dollár (2017), treat object detection as a simple regression problem, straight from image pixels to bounding box coordinates and class probabilities.
 To this end, they produce a dense grid of bounding box proposals and class probabilities in one step.
-Specifically, YOLO divides the input image into a grid and predicts bounding boxes and class probabilities for each grid cell.
-One-stage detectors are usually adopted for real-time applications due to their high speed and efficiency, although their accuracy tends to be lower when being compared with two-stage detectors.
+Specifically, YOLO Terven et al.
+(2023) divides the input image into a grid and predicts bounding boxes and class probabilities for each grid cell.
+Although this efficiency is beneficial for real-time applications, it often comes at the cost of accuracy when compared to two-stage detectors
 Two-stage detectors
@@
 This second stage typically involves a more complex network, such as a convolutional neural network (CNN), which performs classification and further refinement of the bounding box coordinates.
-Two-stage detectors are known for their high accuracy and robustness in detecting objects of varying sizes and shapes, but they are generally slower than one-stage detectors.
+This two-step process boosts accuracy by allowing for a more refined feature analysis, though it also slows down processing, making two-stage detectors less suited for applications that require high-speed performance.
 Most detector networks, including Faster R-CNN and YOLO, produce a large number of bounding box proposals which are subsequently refined using confidence thresholding and Non-Maximum Suppression (NMS) to produce a set of finally detected objects in the image.
@@
 where $\mathbf{L}_{i}$ defines the bounding box corners $(x^{i}_{1},y^{i}_{1})$ and $(x^{i}_{2},y^{i}_{2})$; $O_{i}\in[0,1]$ refers to the probability that bounding box $L_{i}$ contains an object of any class; and $\mathbf{P}_{i}$ is a vector of probabilities $(p^{i}_{1},\ldots,p^{i}_{C})$ representing the probability that region $\mathbf{L}_{i}$ belongs to each of $C$ classes.
-As it can be observed in this above definition, unlike traditional classifiers which assign a single class label to an entire image, object detectors must simultaneously handle both classification and localization.
-This dual task involves predicting not only the class of each object but also its precise location within the image.
-The complexity of this task increases the difficulty of making these models interpretable.
+Unlike traditional classifiers, which assign a single class label to an entire image, object detectors must handle both classification and localization simultaneously.
+This dual task, predicting the class and precise location of each object, increases the complexity of making these models interpretable.
 3.2 Explainable Artificial Intelligence (XAI)
@@
 To this end, multiple XAI techniques have been proposed, which can be classified into four broad categories Ali et al.
-(2023b):
+(2023):
   • Scoop-based techniques focus on the extent of the explanation, providing either local explanations for specific predictions or global explanations for the overall model behavior.
@@
 Therefore, we focus on model-agnostic black-box XAI approaches, which are designed to be architecture-independent, and do not depend at all on the specifics of the model under target.
-Among black-box XAI methods, perturbation-based approaches are commonly used due to their simplicity and effectiveness in revealing which parts of the input are most influential for the model’s predictions.
+Among model-agnostic XAI methods, perturbation-based approaches are commonly used due to their simplicity and effectiveness in revealing which parts of the input are most influential for the model’s predictions.
 Perturbation-based techniques offer a direct way to assess how changes to the input image affect the model’s output.
@@
 4 Materials and Methods
-This section describes the industrial robotics use cases providing the datasets (Subsection 4.1), XAI methods (Subsection 4.3) and the explanation quality metrics (Subsection 4.4) considered in our work.
-The novel XAI technique and quality metrics proposed in this manuscript are also described in detail in Subsection 4.3.
+This section describes the industrial robotics use cases in what refers to the datasets (Section 4.1), object detection model (Section 4.2), XAI methods (Section 4.3) and the explanation quality metrics (Section 4.4) considered in our work.
+The novel XAI technique and quality metrics proposed in this manuscript are also described in Section 4.3.
 4.1 Industrial Robotics Datasets under Consideration
@@
 The first dataset, from PIAP https://piap.lukasiewicz.gov.pl/), involves a collaborative workspace where a human and a robotic arm work together.
-The second dataset, provided by Robotnik https://robotnik.eu/, focuses on a battery assembly area, where a robotic arm assembles components for a battery kit.
+The second dataset, provided by Robotnik https://robotnik.eu/, focuses on a battery assembly area, where a robotic arm assembles components for a battery kit [^2] While the datasets contain a relatively small number of images, this data shortage is typically encountered in real-world industrial scenarios subject to data availability constraints.
+Nevertheless, in the use cases under considerations the contextual and scene variability is minimal, yielding short-tailed distributions of the objects to be detected.
+Therefore, the small datasets described in the paper sufficiently capture the relevant features for the specific object detection tasks addressed by the models..
 Dataset 1: Human-Robot Dataset
@@
 We evaluate four popular methods for generating visual explanations of black-box models: LIME, RISE, D-RISE, and D-MFPP.
-The first two methods, LIME and RISE, were originally developed for image classifiers but can be adapted to object detectors.
-However, they primarily focus on explaining classification aspects and are not capable of addressing localization characteristics.
+The first two methods, LIME and RISE [^4] These XAI methods have been chosen due to their perturbation-based nature, which aligns closely with the methodology followed by the XAI methods D-RISE and D-MFPP proposed in this work.
+Both D-RISE and D-MFPP generate explanations through perturbations., were originally developed for image classifiers but can be adapted to object detectors, However, they primarily focus on explaining classification aspects and are not capable of addressing localization characteristics.
 In contrast, D-RISE is one of the first XAI methods specifically designed for object detectors, providing explanations that encompass both classification and localization.
@@
 (2023), explanation quality metrics can be grouped into six categories based on their logical similarity: faithfulness, robustness, localization, complexity, randomization, and axiomatic metrics.
-In this study, we focus on two of these categories that are particularly relevant to object detection: localization (Subsection 4.4.1) and faithfulness (Subsection 4.4.2).
+In this study, we focus on two of these categories that are particularly relevant to object detection: localization (Section 4.4.1) and faithfulness (Section 4.4.2).
 4.4.1 Localization
@@
 Given the importance score for each pixel ($S_{I,f}$), Min-Subset is defined as the smallest subset of pixels that needs to be removed to change the model’s prediction.
-Mathematically,
+Mathematically:
 […]
@@
 This figure highlights how D-Deletion distinguishes between different objects of the same class by incorporating localization information, leading to more refined and accurate explanations ($\downarrow$ AUC in the Figure’s last row) when multiple objects of the same class are detected in an image.
-For the sake of clarity, we provide the pseudocode of Deletion in Algorithm 1, where the main difference with respect to D-Deletion are lines 9 to 11.
+For the sake of clarity, we provide the pseudocode of Deletion in Algorithm 1, where the main difference with respect to D-Deletion are lines 10 to 12.
 Lastly, akin to how Min-Subset is related to Deletion, D-Min-Subset is associated with D-Deletion.
@@
 In this context, to evaluate the effectiveness of explanations, we formulate four key research questions to answer them with empirical evidence:
-  • RQ1. Which XAI method provides the most reliable and insightful explanations for object detection models?
-  • RQ2. Does the D-Deletion metric enhance the trustworthiness of XAI outputs when multiple objects of the same class are present in the image?
-  • RQ3. How does the mask generation process influence the quality of explanations, particularly when using similarity scores for object detection? How does D-MFPP behave?
-  • RQ4.
-Do different image dimensions impact the explanations generated by XAI methods?
+  • RQ1: Which XAI method provides the most reliable and insightful explanations for object detection models?
+  • RQ2: Does the D-Deletion metric enhance the trustworthiness of XAI outputs when multiple objects of the same class are present in the image?
+  • RQ3: How does the mask generation process influence the quality of explanations, particularly when using similarity scores for object detection? How does D-MFPP behave?
+  • RQ4: Do different image dimensions impact the explanations generated by XAI methods?
 Do models of varying sizes (large, medium, small, nano) focus on different regions of the image in their explanations?
@@
 In what follows we present and discuss on the results obtained to answer each of the RQ formulated above:
-RQ1. Comparison between XAI methods
+RQ1: Comparison between XAI methods
 Figure 5: Heatmaps obtained by applying RISE (left) and D-RISE (right) for the detection of a human in the Human-Robot Dataset.
-Figure 6: Heatmaps generated in a scene of the Battery Assembly dataset for two target classes: individual battery (top row) and battery holder (bottom row).
-The first and second columns shows the saliency maps obtained using LIME and RISE, independent of the number of objects of the same class in the image.
-Columns 3,4 and 5 display heatmaps generated using D-RISE for three different individual elements of the same class.
+Figure 6: Explanations of a scene using different stride and window size configurations when using D-Sliding Window (combination of mask generation explained in Section 3.2 and Equation (6)).
+(Left) Stride of 16; (Right) Stride of 8; (Top) Window size of 32; (Bottom) Window size of 64.
+Table 1: Quantitative metrics of LIME, RISE and D-RISE for the Human-Robot dataset.
+The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each class (Human, Gripper) and the overall average.
+Lower values are better for metrics marked with $\downarrow$, while higher values are better for those marked with $\uparrow$.
+Bold values indicate the best average scores across all objects, highlighting the best-performing XAI method for each metric.
+Values highlighted in gray represent the best scores for each object category (Human or Gripper) and should be interpreted vertically.
+XAI Method Object Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
+LIME Human 0.0759 0.0632 4.2703 4.2703 1.0000 34.984
+Gripper 0.4688 0.0324 1.3859 1.3859 1.0000 2.2270
+Average 0.2723 0.0478 2.8281 2.8281 1.0000 18.6060
+RISE Human 0.1827 0.1241 9.0108 9.0108 0.7500 19.0824
+Gripper 0.2637 0.0060 0.6355 0.63 1.0000 1.0542
+Average 0.2232 0.0651 4.8232 4.8232 0.875 10.0683
+D-RISE Human 0.1255 0.0818 5.7335 5.7335 0.8750 20.4061
+Gripper 0.2777 0.0059 0.6091 0.6091 1.0000 1.0815
+Average 0.2016 0.0438 3.1713 3.1713 0.9375 10.7438
 In the Human-Robot dataset, the comparison between LIME, RISE, and D-RISE, as shown in Table 1, reveals distinct strengths across different metrics (Section 4.4).
@@
 However, D-RISE eliminates less relevant areas more effectively.
-Table 1: Quantitative metrics of LIME, RISE and D-RISE for the Human-Robot dataset.
-The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each class and the overall average.
-$\downarrow$ indicates that lower values are better, whereas $\uparrow$ indicate that higher values are better.
-The best results for each metric and object type are highlighted in gray, while the best average values are written in bold.
-XAI Method Object Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
-LIME Human 0.0759 0.0632 4.2703 4.2703 1.0000 34.984
-Gripper 0.4688 0.0324 1.3859 1.3859 1.0000 2.2270
-Average 0.2723 0.0478 2.8281 2.8281 1.0000 18.6060
-RISE Human 0.1827 0.1241 9.0108 9.0108 0.7500 19.0824
-Gripper 0.2637 0.0060 0.6355 0.63 1.0000 1.0542
-Average 0.2232 0.0651 4.8232 4.8232 0.875 10.0683
-D-RISE Human 0.1255 0.0818 5.7335 5.7335 0.8750 20.4061
-Gripper 0.2777 0.0059 0.6091 0.6091 1.0000 1.0815
-Average 0.2016 0.0438 3.1713 3.1713 0.9375 10.7438
 In the results obtained over the Battery Assembly dataset (Table 2), a similar pattern can be noticed.
@@
 D-RISE also shows the highest PG score (97%), performing significantly better than LIME (76.85%) and RISE (66.95%).
+Overall, when dealing with datasets containing only one object per class, the differences between LIME, RISE, and D-RISE are relatively small in quantitative terms.
+However, when multiple objects of the same class appear in a given input image, D-RISE clearly dominates over the rest of techniques.
+As illustrated in Figure 7, D-RISE generates coherent heatmaps for each detected object in the Battery Assembly dataset, whereas LIME and RISE provide a global saliency map for the entire class.
+By combining the individual saliency maps from D-RISE, a more accurate and object-specific explanation can be produced.
+This also highlights the limitations of LIME and RISE when applied to multiple objects, as their global saliency maps do not differentiate between individual instances.
 Table 2: Quantitative metrics of LIME, RISE and D-RISE for the Battery Assembly dataset.
-The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each object and the overall average.
-The best results for each metric and object type are highlighted in gray, while the best average values are marked in bold.
+The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each object (indiv batt, bms a, bms b, unknown object, batt holder) and the overall average.
+Lower values are better for metrics marked with $\downarrow$, while higher values are better for those marked with $\uparrow$.
+Bold values indicate the best average scores across all objects, highlighting the best-performing XAI method for each metric.
+Gray-highlighted values represent the best scores for each object category (indiv batt, bms a, bms b, unknown object or batt holder) and should be interpreted vertically.
 XAI Method Object Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
@@
 Average 0.2448 0.0344 12.9141 3.0819 0.9714 4.1768
-Overall, when dealing with datasets containing only one object per class, the differences between LIME, RISE, and D-RISE are relatively small in quantitative terms.
-However, when multiple objects of the same class appear in a given input image, D-RISE clearly dominates over the rest of techniques.
-As illustrated in Figure 6, D-RISE generates coherent heatmaps for each detected object in the Battery Assembly dataset, whereas LIME and RISE provide a global saliency map for the entire class.
-By combining the individual saliency maps from D-RISE, a more accurate and object-specific explanation can be produced.
-This also highlights the limitations of LIME and RISE when applied to multiple objects, as their global saliency maps do not differentiate between individual instances.
+Figure 7: Heatmaps generated in a scene of the Battery Assembly dataset for two target classes: individual battery (top row) and battery holder (bottom row).
+The first and second columns shows the saliency maps obtained using LIME and RISE, independent of the number of objects of the same class in the image.
+Columns 3,4 and 5 display heatmaps generated using D-RISE for three different individual elements of the same class.
+RQ2: D-Deletion metric for scenes with multiple objects of the same class
+As a secondary observation in the experiments of RQ1, the D-Deletion metric is specifically designed to overcome the limitations of traditional deletion metrics, particularly when multiple objects of the same class are present in an image.
+In the Battery Assembly dataset, where several instances of the same class (e.g., indiv batt) appear, D-Deletion demonstrates clear advantages.
+By inspecting Table 2, RISE, while performing reasonably well with an average Deletion score of $0.2860$, it still obtains a relatively high D-Deletion score of 0.1474, suggesting that it struggles to differentiate between the contributions of individual objects.
+In contrast, D-RISE, which obtains an average Deletion score of 0.2448, outperforms RISE with a D-Deletion score of 0.03444.
+This highlights D-RISE’s ability to isolate and preserve key features for each object, providing more trustworthy, object-specific explanations rather than broad, class-level insights.
+The Min-Subset and D-Min-Subset metrics, which measure the minimal proportion of pixels needed to disrupt a detection, reinforce these findings.
+In the Human-Robot dataset, Table 1, where only one object per class appears, the differences between Deletion and D-Deletion scores are minor, and the Min- Subset and D-Min-Subset values are close to each other.
+However, in the Battery Assembly dataset, where the differences between Deletion and D-Deletion are more substantial and multiple objects of the same class co-occur in the same image, the Min-Subset ($12.9141$) and D-Min-Subset ($3.0819$) values also diverge significantly.
 Class XAI $N=500$ $N=1,000$ $N=5,000$ $N=10,000$
@@
 Saliency maps are presented for two scenes, one from each dataset, demonstrating the evolution of heatmaps for a single element as the number of masks increases, highlighting the effect on explanation quality and focus.
-RQ2. D-Deletion metric for scenes with multiple objects of the same class
-As a secondary observation in the experiments of RQ1, the D-Deletion metric is specifically designed to overcome the limitations of traditional deletion metrics, particularly when multiple objects of the same class are present in an image.
-In the Battery Assembly dataset, where several instances of the same class (e.g., indiv batt) appear, D-Deletion demonstrates clear advantages.
-By inspecting Table 2, RISE, while performing reasonably well with an average Deletion score of $0.2860$, it still obtains a relatively high D-Deletion score of 0.1474, suggesting that it struggles to differentiate between the contributions of individual objects.
-In contrast, D-RISE, which obtains an average Deletion score of 0.2448, outperforms RISE with a D-Deletion score of 0.03444.
-This highlights D-RISE’s ability to isolate and preserve key features for each object, providing more trustworthy, object-specific explanations rather than broad, class-level insights.
-The Min-Subset and D-Min-Subset metrics, which measure the minimal proportion of pixels needed to disrupt a detection, reinforce these findings.
-In the Human-Robot dataset, Table 1, where only one object per class appears, the differences between Deletion and D-Deletion scores are minor, and the Min- Subset and D-Min-Subset values are close to each other.
-However, in the Battery Assembly dataset, where the differences between Deletion and D-Deletion are more substantial and multiple objects of the same class co-occur in the same image, the Min-Subset ($12.9141$) and D-Min-Subset ($3.0819$) values also diverge significantly.
-RQ3. Influence of the Mask Generation Strategy
-Figure 8: Explanations of a scene using different stride and window size configurations when using D-Sliding Window (combination of mask generation explained in Section 3.2 and Equation (6)).
-(Left) Stride of 16; (Right) Stride of 8; (Top) Window size of 32; (Bottom) Window size of 64.
+Table 3: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Human-Robot dataset.
+Only the average score across all classes is presented.
+Results are interpreted vertically within each column (each metric represents a separate column).
+The best mask configuration for each XAI method and for each metric is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
+.
+XAI Method Num Masks Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
+D-Sliding Window (*) $w=32$, $s=8$ 0.2323 0.2323 24.1966 24.1966 1.0000 97.7128
+$w=32$, $s=16$ 0.3009 0.3009 34.3990 34.3990 1.0000 96.5487
+$w=64$, $s=8$ 0.2046 0.2002 20.7980 20.7980 1.0000 83.9141
+$w=64$, $s=16$ 0.2039 0.1995 20.7362 20.7362 1.0000 83.5582
+D-RISE 500 0.2645 0.0702 4.3520 4.3520 0.9062 10.7117
+1000 0.2530 0.0519 4.2019 4.2019 0.9375 10.8382
+5000 0.2016 0.0438 3.1713 3.1713 0.9375 10.7438
+10000 0.2246 0.0389 3.2496 3.2496 0.9375 10.7094
+D-MFPP 500 0.4137 0.0505 2.9608 2.9519 1.0000 15.9700
+1000 0.4227 0.0678 5.8648 5.8648 1.0000 15.1166
+5000 0.3944 0.0469 3.8146 3.8058 1.0000 16.0809
+10000 0.3893 0.0471 3.6734 3.6734 1.0000 16.2579
+*: D-Sliding Window scores have been computed excluding the cases where no explanations are produced for images with large-sized object instances.
+Table 4: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Battery Assembly dataset.
+Results are interpreted vertically within each column (each metric represents a separate column).
+The best mask configuration for each XAI method and for each metric is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
+XAI Method Num Masks Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
+D-Sliding Window (*) $w=32$, $s=8$ 0.5198 0.3258 53.9463 31.2219 0.9855 90.4736
+$w=32$, $s=16$ 0.4700 0.3046 45.9925 24.7660 0.9788 84.5302
+$w=64$, $s=8$ 0.2578 0.0245 26.4318 2.4062 0.9855 55.8151
+$w=64$, $s=16$ 0.2570 0.0246 26.3537 2.4772 0.9884 54.8283
+D-RISE 500 0.4307 0.0783 29.7843 4.9001 0.9971 3.8152
+1000 0.3837 0.0701 26.4242 3.3443 0.9971 3.8154
+5000 0.3645 0.0431 25.2766 3.1374 0.9971 3.8015
+10000 0.3707 0.0342 23.3216 2.9789 0.9971 3.7883
+D-MFPP 500 0.3327 0.0708 30.7555 5.6594 0.9942 6.2320
+1000 0.3375 0.0630 30.0871 4.8843 0.9971 6.1938
+5000 0.3042 0.0427 27.8461 3.4328 0.9971 6.1772
+10000 0.3100 0.0409 29.6047 3.8185 0.9971 6.1680
+*: D-Sliding Window scores have been computed excluding the cases where no explanations are produced for images with large-sized object instances.
+RQ3: Influence of the Mask Generation Strategy
 When comparing XAI approaches for object detection tasks configured with different mask generation techniques, the results in Tables 3 and 4 initially suggest that D-Sliding Window performs the best in almost all metrics.
@@
 This trend can be also noted in Figure 8, where (1) increasing the number of masks improves the quality of the saliency maps, and (2) D-MFPP provides cleaner maps with a lower number of masks.
-Table 3: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Human-Robot dataset.
-Only the average score across all classes is presented.
-The best mask configuration for each XAI method is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
-XAI Method Num Masks Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
-D-Sliding Window (*) $w=32$, $s=8$ 0.2323 0.2323 24.1966 24.1966 1.0000 97.7128
-$w=32$, $s=16$ 0.3009 0.3009 34.3990 34.3990 1.0000 96.5487
-$w=64$, $s=8$ 0.2046 0.2002 20.7980 20.7980 1.0000 83.9141
-$w=64$, $s=16$ 0.2039 0.1995 20.7362 20.7362 1.0000 83.5582
-D-RISE 500 0.2645 0.0702 4.3520 4.3520 0.9062 10.7117
-1000 0.2530 0.0519 4.2019 4.2019 0.9375 10.8382
-5000 0.2016 0.0438 3.1713 3.1713 0.9375 10.7438
-10000 0.2246 0.0389 3.2496 3.2496 0.9375 10.7094
-D-MFPP 500 0.4137 0.0505 2.9608 2.9519 1.0000 15.9700
-1000 0.4227 0.0678 5.8648 5.8648 1.0000 15.1166
-5000 0.3944 0.0469 3.8146 3.8058 1.0000 16.0809
-10000 0.3893 0.0471 3.6734 3.6734 1.0000 16.2579
-*: D-Sliding Window scores have been computed excluding the cases where no explanations are produced for images with large-sized object instances.
-Table 4: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Battery Assembly dataset.
-The best mask configuration for each XAI method is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
-XAI Method Num Masks Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
-D-Sliding Window (*) $w=32$, $s=8$ 0.5198 0.3258 53.9463 31.2219 0.9855 90.4736
-$w=32$, $s=16$ 0.4700 0.3046 45.9925 24.7660 0.9788 84.5302
-$w=64$, $s=8$ 0.2578 0.0245 26.4318 2.4062 0.9855 55.8151
-$w=64$, $s=16$ 0.2570 0.0246 26.3537 2.4772 0.9884 54.8283
-D-RISE 500 0.4307 0.0783 29.7843 4.9001 0.9971 3.8152
-1000 0.3837 0.0701 26.4242 3.3443 0.9971 3.8154
-5000 0.3645 0.0431 25.2766 3.1374 0.9971 3.8015
-10000 0.3707 0.0342 23.3216 2.9789 0.9971 3.7883
-D-MFPP 500 0.3327 0.0708 30.7555 5.6594 0.9942 6.2320
-1000 0.3375 0.0630 30.0871 4.8843 0.9971 6.1938
-5000 0.3042 0.0427 27.8461 3.4328 0.9971 6.1772
-10000 0.3100 0.0409 29.6047 3.8185 0.9971 6.1680
-*: D-Sliding Window scores have been computed excluding the cases where no explanations are produced for images with large-sized object instances.
-RQ4. Image dimension variation and YOLOv8 complexity
-Table 5: Explanations comparison for three image dimensions (720x1280, 360x640, and 180x320) and varying resolutions (16x16, 8x8, and 4x4) in the Human-Robot dataset.
+RQ4: Image dimension variation and YOLOv8 complexity
+Table 5: Explanations comparison for three image dimensions ($720\times 1280$, $360\times 640$, and $180\times 320$) and varying resolutions ($16\times 16$, $8\times 8$, and $4\times 4$) in the Human-Robot dataset.
 Image dimensions Resolution Deletion ($\downarrow$) D-Deletion ($\downarrow$) Min-Subset ($\downarrow$) D-Min-Subset ($\downarrow$) PG ($\uparrow$) EBPG ($\uparrow$)
@@
 Laña also acknowledge funding from the Basque Government (MATHMODE, IT1456-22).
+Author Contributions
+Alain Andres: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Writing (Original draft preparation).
+Aitor Martinez-Seras: Validation, Writing (Original draft preparation).
+Ibai Laña: Supervision, Project administration, Writing (Review and Editing).
+Javier Del Ser: Conceptualization, Supervision, Validation, Writing (Reviewing and Editing).
 Declaration of competing interest
 The authors declare that they have no conflicts of interest regarding this work.
-Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union.
+However, views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union.
 The European Union can not be held responsible for them.